### v1

# Improving absolute separability bounds for arbitrary dimensions

Abstract
Sufficient analytical conditions for separability in composite quantum systems are very scarce and only known for low-dimensional cases. Here, we use linear maps and their inverses to derive powerful analytical conditions, providing tight bounds and extremal points of the set of absolutely separable states, i.e., states that remain separable under any global unitary transformation. Our analytical results apply to generic quantum states in arbitrary dimensions, and depend only on a single or very few eigenvalues of the considered state. Furthermore, we use convex geometry tools to improve the general characterization of the AS set given several non-comparable criteria. Finally, we present various conditions related to the twin problem of characterizing absolute PPT, that is, the set of quantum states that are positive under partial transposition and remain so under all unitary transformations.
footnotetext: *These authors contributed equally to this work.
1 Introduction
As it is well known, determining and quantifying entanglement has been proven to be, in general, an NP-Hard problem [1, 2]. After several decades of intense research, sufficient and necessary conditions to ensure or deny the entanglement content of a state exist only for the simplest cases [3, 4], or very restricted families of states. One of those conditions is linked to positive but not completely positive maps, but to date neither the use of a finite number of such maps nor the use of semidefinite programming computation [5, 6] can give a conclusive answer if a given state is entangled or not [7, 8]. Under these circumstances, one may relax the quest for iff conditions and search for sufficient conditions for separability.
Here, we wish to draw attention to the use of not completely positive maps and their inverses, as sufficient separability criteria [9, 10, 11]. In particular, we focus here on the concept of absolute separability (AS), that is, those states whose separability is preserved under all unitary transformations. Characterizing the set of AS is relevant for various reasons. Asides of its importance for the general characterization of the set of quantum states, AS states become pertinent in the resource theory of entanglement, as they represent free states, while entangled states act as a resource. Since unitary transformations are the most common method to generate entanglement, e.g., through quantum circuits or quenches, if a state is certified AS it cannot lead to any entanglement under unitary transformations [12]. Moreover, AS states depend only on their spectrum (eigenvalues), and their characterization is significantly less demanding than performing full quantum state tomography [13, 14]. For a quantum state of dimension $D=N\cdot M$ (with $N,M$ representing a bipartition), full tomography requires reconstructing $D^{2}-1$ real parameters, while only $D-1$ eigenvalues are needed to characterize the spectrum. Shadow tomography [15], which is restricted to the possible outcomes of the observables to be measured, also scales less favorable than $D-1$. Finally, we note that the AS set is convex, and thus its characterization relies on determining its extremal points.
Several sufficient criteria for AS have been provided over the years. For instance, by bounding matrix norms [16, 17], through linear matrix inequalities [18, 19], by proving the equivalency of absolute separability and absolute PPT for certain subsets [20, 21], using separability witnesses [22] or searching for extremal points [23] among others. Despite these previous achievements, to this date, the set of AS states remains not fully characterized [24].
In this work, we provide a finer geometric characterization of the AS set for arbitrary dimensions. Our results lead to compact analytical conditions of AS which rely on very few or even a single eigenvalue of the state. With similar techniques, we also provide tighter bounds on the geometry of AS set in the multipartite scenario, both for qubits or qudits.
Due to the convexity of the AS set, criteria for AS may be expressed as convex conditions on the state $\rho$. This observation reflects that quantum entanglement, as a resource, cannot be produced (nor amplified) by mere statistical mixing of different separable states. One goal of the present contribution is to join previous criteria with ours into a unifying framework based on convex programs, a class of efficiently solvable optimization algorithms. The separability tests proposed here reconcile the different separability conditions, but also enable to systematically improve over them, for free, without requiring additional experimental effort. In doing so, we bring a new perspective on a long-standing open problem in quantum information theory [8].
The paper is structured as follows: in Section 2, we introduce the formal definitions of AS and its twin problem of absolute positive partial transposition (AP). We review previous criteria, and derive new theorems for AS based on inverting non-completely positive maps. In Section 3, we present our methods to improve present bounds of absolute separability using convex geometrical tools. Importantly, our methods apply to arbitrary dimensions. In Section 4, we address absolute separability in multipartite systems where, again, we derive conditions based on a single eigenvalue alone. Finally, we focus on the symmetric subspace, where analogous conditions for AS and AP can be established. We conclude in Section 5.
2 Absolute separability
A quantum state is separable if it can be represented as a convex combination of product states. For bipartite states, $\rho\in\mathcal{B}(\mathbb{C}^{N}\otimes\mathbb{C}^{M})$, (where $\mathcal{B}(\mathcal{H})$ denotes the set of bounded operators acting on $\mathcal{H}$), that implies
$\rho=\sum_{i}p_{i}\ket{e_{i},f_{i}}\bra{e_{i},f_{i}},$
[…]
Definition 1.
A state $\rho\in\mathcal{B}(\mathbb{C}^{N}\otimes\mathbb{C}^{M})$ is absolutely separable (AS) if and only if, for any unitary matrix $U$ acting on $\mathbb{C}^{N}\otimes\mathbb{C}^{M}$, $\rho^{\prime}=U\rho U^{{\dagger}}$ is separable.
Since unitary transformations preserve the spectrum of a given density matrix $\rho$, it is possible to infer absolute separability criteria based solely on the eigenvalues of the state [20]. For that reason, the set AS is also called separable from spectrum. Clearly, the maximally mixed state (MMS), $\rho\propto\mathds{1}$, is absolutely separable, hence the AS set must contain the MMS and states sufficiently close to it.
The most celebrated criteria for AS is based in the matrix norm of the operator $\rho$, and it is usually referred to as the Gurvits-Barnum ball (GBB) [16]. This criterion can be expressed as a function of the purity of the state, $\mathrm{Tr}(\rho^{2})$, as presented in the following theorem.
Theorem 1.
[16] Let $\rho\in\mathcal{B}(\mathbb{C}^{N}\otimes\mathbb{C}^{M})$ be a normalized state in a Hilbert space of global dimension $D=N\cdot M$. If
$\mathrm{Tr}(\rho^{2})\leq\frac{1}{D-1},$ (1)
then $\rho$ is absolutely separable.
From the spectral decomposition of the state $\rho=\sum_{i=0}^{N\cdot M-1}\lambda_{i}\outerproduct{\psi_{i}}{\psi_{i}}$ |, the condition given by Eq. (1) is geometrically interpreted as a ball of radius $(\sqrt{D-1})^{-1}$ in the space of eigenvalues $\bm{\lambda}=\{\lambda_{i}\}$. For normalized states, i.e. $\mathrm{Tr}(\rho)=\sum_{i=0}^{NM-1}\lambda_{i}=1$, this ball is centered on the maximally mixed state.
Absolute separability from linear maps.– Here, we take a different approach, based on linear maps, to tackle the characterization of AS. We recall that positive but not completely positive linear maps, $\Lambda:\mathcal{B}(\mathcal{H})\rightarrow\mathcal{B}(\mathcal{H}^{\prime})$, provide sufficient conditions for entanglement [4, 25, 26] (or, equivalently, necessary conditions for separability) in the form of entanglement witnesses. Interestingly, as reported in [9, 11], such maps are also instrumental to derive sufficient conditions for separability. Before proceeding further, we present the central result on which our work relies on:
Theorem 2.
[9] Let S, S’ be convex and compact subsets of the space of all bounded operators $\mathcal{B}(\mathbb{C}^{N}\otimes\mathbb{C}^{M})$, and let $\Lambda_{\mathbf{p}}:S\rightarrow S^{\prime}$ be a family of maps, invertible for almost all $\mathbf{p}$. By $\mathcal{P}_{SS^{\prime}}$ we denote the subset of the parameters set $\mathbf{p}$, for which the maps have the property that, for every $\rho\in S$, $\Lambda_{\mathbf{p}}(\rho)\in S^{\prime}$ provided $\mathbf{p}\in\mathcal{P}_{SS^{\prime}}$. Then, if $\Lambda_{\mathbf{p}}^{-1}(\sigma)\in S$, it follows that $\sigma\in S^{\prime}$.
In other words, if the set S corresponds to all quantum states acting on $\mathbb{C}^{N}\otimes\mathbb{C}^{M}$, and the set $S^{\prime}$ is the subset of separable states, the above theorem provides a sufficient criterion for certifying that a a given state $\sigma\in\mathcal{B}(\mathbb{C}^{N}\otimes\mathbb{C}^{M}$) is separable. As pointed out in [9], the choice of $\Lambda_{\mathbf{p}}$ is determined first by the capability of proving the assumption that $\Lambda_{\bf p}(S)\subset S^{\prime}$, and second by checking that $\Lambda_{\bf p}^{-1}(S^{\prime})\subset S$, that is, by the difficulty of inverting analytically the chosen map.
In this work, we aim at deriving sufficient criteria to detect AS from the previous theorem. To that end, we consider $S$ to be the set of all quantum states, i.e., $\rho\geq 0$ (or, up to normalization, positive-semi definite (PSD) matrices), and $S^{\prime}$ a subset of states which are separable from their spectrum. For this endeavor, it is natural to consider unitarily equivariant linear maps $\Lambda$, i.e., those fulfilling that $\forall\rho,U$ acting on $\mathbb{C}^{N}\otimes\mathbb{C}^{M}$, $\Lambda(U\rho U^{{\dagger}})=U\Lambda(\rho)U^{{\dagger}}$. In this direction, it has been shown [27, 28] that the only unitarily equivariant linear maps are the so-called reduction-like maps [29], which are parameterized as:
[…]
with the single parameter $\alpha\in\mathbb{R}$. This family of linear maps is invertible for any nonzero $\alpha$. Its inverse, which is obviously linear, reads:
[…]
where $D$ is the dimension of the Hilbert space in which $\sigma$ acts.
As demonstrated in [9], the map, $\Lambda_{\alpha}(\rho)$, transforms any state $\rho$ into a separable state for $\alpha\in[-1,2]$. As a result, Theorem 2 shows that the positivity of the inverted map, $\Lambda_{\alpha}^{-1}(\sigma)$ serves as a sufficient condition for separability. By noting that the positivity of $\Lambda_{\alpha}^{-1}(\sigma)$ depends only on the spectrum of state $\sigma$, and that the map is unitarily invariant, it immediately becomes a sufficient condition to certify AS. Moreover, as the conditions for positivity are bounded by the values $\alpha=-1$, and $\alpha=2$, these values lead to sufficient conditions for AS from just a single eigenvalue (either the minimum or maximum).
Theorem 3.
Let $\rho$ be a normalized bipartite state acting on the space $\mathbb{C}^{N}\otimes\mathbb{C}^{M}$ with minimal and maximal eigenvalues $\lambda_{\min}(\rho)$ and $\lambda_{\max}(\rho)$ respectively. If
$\lambda_{\min}(\rho)\geq\frac{1}{N\cdot M+2}\quad\text{or}\quad\lambda_{\max}(\rho)\leq\frac{1}{N\cdot M-1},$ (4)
then $\rho$ is absolutely separable.
Proof.
(See also [9, 30]). For $02\geq\alpha>$ 0, positivity of $\Lambda_{\alpha}^{-1}(\rho)$ demands that $\rho-\mathds{1}/(N\cdot M+\alpha)\geq 0$, that is, $\lambda_{\min}(\rho)-1/(M\cdot N+\alpha)\geq 0$. On the other hand, for $10>\alpha\geq-$ 1, the condition reads $\mathds{1}/(M\cdot N+\alpha)-\rho\geq 0$, or equivalently, $1/(M\cdot N+\alpha)-\lambda_{\max}(\rho)\geq 0$. The extreme values $\alpha=2$ and $\alpha=-1$ yield the announced result. ∎
Finally, let us remark that both, the conditions contained in Theorem 3 and the conditions on Gurvits-Barnum ball, Theorem 1, depend solely on the dimension of the global bipartite system $D=N\cdot M$, independently of the specific local dimensions $N$ and $M$. For a multipartite state of global dimension $D$, they provide sufficient conditions for absolute separability with respect to any bipartition of the total system (see Section 4). As we will show later, the conditions derived from Theorem 3, moreover, provide extremal points in the convex set of the AS set, and permit a geometrical approach to the set of AS states. The characterization of the extreme points of the AS states has been recently achieved but only for specific dimensions; the qubit-qudit and qutrit-qudit case [21, 23].
Recall that all separable states are necessarily positive under partial transposition (PPT), as stated by the famous Peres-Horodecki criterion [3, 31]. It is a natural to ask if there is an analogous set of states that remain positive under partial transposition under the action of any unitary transformations (AP), and if such a set coincides with the set of absolute separable states (AS) [19, 24]. In analogy to the set of AS, the set AP is defined as:
Definition 2.
A normalized state $\rho\in\mathcal{B}(\mathbb{C}^{N}\otimes\mathbb{C}^{M})$ is absolutely positive under partial transposition (AP) if and only if, for any unitary matrix $U$ acting on $\mathbb{C}^{N}\otimes\mathbb{C}^{M}$, $\rho^{\prime}=U\rho U^{{\dagger}}$ is positive under partial transposition.
In contrast to the AS set, the set of AP has been completely characterized in arbitrary dimensions by a series of linear matrix inequalities (LMI) derived originally in [18]. However, the number of LMI that is necessary to certify AP grows exponentially with $\min\{N,M\}$, which prohibits the characterization for high dimensions. A necessary condition for AP can be derived from the positivity of the first $2\times 2$ block of the first LMI (see Theorem 1 and Section 4 of [18]). Then, considering the eigenvalues of a state $\rho\in\mathcal{B}(\mathbb{C}^{N}\otimes\mathbb{C}^{M})$ in non-decreasing order, that is, $\lambda_{0}\leq\lambda_{1}\leq\cdots\leq\lambda_{N\cdot M-1}$, a necessary condition for AP reads as
$\begin{pmatrix}2\cdot\lambda_{0}&\lambda_{1}-\lambda_{NM-1}\\ \lambda_{1}-\lambda_{NM-1}&2\cdot\lambda_{2}\end{pmatrix}\geq 0,$ (5)
which involves only the three lowest eigenvalues, $\{\lambda_{0}\leq\lambda_{1}\leq\lambda_{2}\}$, and the largest one $\lambda_{N\cdot M-1}$. From Eq. (5) and Theorem 3, and considering the vectors of eigenvalues given by $(1-\frac{D-1}{D+\alpha},\frac{1}{D+\alpha},\cdots,\frac{1}{D+\alpha})$, it is possible to check that only for $\alpha\in[-1,2]$ the necessary condition for AP is fulfilled. Thus, the region for AP coincides with the one for AS.
Remarkably, only for the qubit-qudit case it has been proven that PPT is also sufficient for AS [21]. For arbitrary bipartitions, the equivalence between the AS and AP sets is only conjectured [24]. Although the proposed Theorem 3 does not provide a complete characterization of the AS set, it is valid for any local dimensions and supports the conjecture that AS and AP are identical sets.
3 Improving sufficient criteria for absolute separability in arbitrary dimensions from convex geometry
Having set the main tools of our analysis, here we show how to improve present bounds of absolute separability based on convex geometrical methods. The set of states detected by the inverse of the reduction map (Theorem 3) for different values of $\alpha$ and the states enclosed in Gurvits-Barnum ball (Theorem 1) are not comparable. Nevertheless, since the set of AS states is convex and compact [32], merging both criteria with extreme values of $\alpha_{\pm}$ into a single convex hull can yield stronger separability conditions.
3.1 Detecting separability from few eigenvalues
The subsets of AS states in $\mathbb{C}^{N}\otimes\mathbb{C}^{M}$ detected by Theorem 3, form simplexes in the space of eigenvalues. For the case of arbitrary dimensions $D=N\cdot M$, the vertices of these simplexes correspond to all permutations on the vector formed by the eigenvalues. Explicitly, for $\alpha=2$, the largest simplex corresponds to the vector $\left(\frac{1}{D+2},\cdots,\frac{1}{D+2},1-\frac{D-1}{D+2}\right)$ and its permutations, while for $\alpha=-1$, the vertices are settled by $\left(0,\frac{1}{D-1},\cdots,\frac{1}{D-1}\right)$ and its permutations. A schematic representation of these sets for $D=3$ is depicted in Figure 1 using barycentric coordinates $\{\tilde{\lambda}_{i}\}$, where the maximally mixed state (MMS) lies at the origin and pure states at the vertexes of the simplex of normalized states.
$\tilde{\lambda}_{1}$$\tilde{\lambda}_{2}$
Figure 1: Schematic representation of the set of states and absolutely separable states (AS) for the case of $D=3$ in barycentric coordinates. The normalized set of states is represented by the area enclosed in the blue vertices. The two simplexes obtained from Theorem 3, corresponding to the conditions $\lambda_{\min}(\rho)\geq 1/(D+2)$ (red dashed line) and the set with $\lambda_{\max}(\rho)\leq 1/(D-1)$ (orange dotted line), are also plotted. The thick green line displays the convex hull of both simplexes. Finally, the shaded zone fulfills the ordering $\lambda_{0}\leq\lambda_{1}\leq\lambda_{2}$ that we consider across this work without loss of generality. Note that the case $D=3$ does not correspond to a physical system, as it cannot factorize into any subsystems; however, it is depicted here to exemplify the general rule.
The displayed simplexes of AS are, in fact, extreme due to the tightness of the values of $\alpha$ displayed in Eq. (3). Thus, the convex hull arising from the union of both simplexes will also be optimal, i.e., the biggest possible one. Through convexity of the AS set, it is also possible to upper bound it, since the vertex of a simplex with $\alpha\pm\varepsilon$ will not be AS. We formally state these facts in the following proposition.
Proposition 1.
Let $\rho$ be a bipartite state acting in a Hilbert space $\mathbb{C}^{N}\otimes\mathbb{C}^{M}$ of dimension $D=N\cdot M$, and let $\{\lambda_{i}\}_{i=0}^{D-1}$ denote its corresponding eigenvalues in non-decreasing order $\{\lambda_{i}\leq\lambda_{i+1}\}_{i=0,1,..,D-1}$. If $\rho$ is contained in the CH of the conditions presented in Theorem 3, then
$3\cdot\sum_{i=0}^{\lfloor\frac{D+1}{3}\rfloor-1}\lambda_{i}+\left(D+2-3\cdot\left\lfloor\frac{D+1}{3}\right\rfloor\right)\cdot\lambda_{\lfloor\frac{D+1}{3}\rfloor}\geq 1$ (6)
Proof.
In Appendix A, we show how to numerically obtain the proposed solution as an ansatz given a specific dimension. In Appendix B, by assessing the Schur-convexity of the condition and majorization over the vectors of eigenvalues, we prove the statement. ∎
For low dimensional systems ($D\leq 16$), we verify that the converse relation also holds. Moreover, for higher dimensions, we obtain numerical evidence that only points contained in the CH are detected by the single inequality. Thus, Eq. 6 yields a proper criterion for AS.
The condition provided by Eq. (6) holds for highly mixed states, close to the MMS, with the specified combination of the smallest eigenvalues large enough. Provided the ordering, $\{\lambda_{i+1}\geq\lambda_{i}\}$, one can replace $\lambda_{i+1}$ by $\lambda_{i}$ in Eq. (6) to yield new sufficient conditions for AS. Albeit being weaker, such conditions are based on fewer eigenvalues than the original theorem, which enables the detection of AS from incomplete knowledge of the spectrum. By iteratively applying the above substitution for $i\leq\kappa\leq\lfloor\frac{D+1}{3}\rfloor$, we derive a hierarchy of AS conditions as follows:
$\kappa=0:$ $\quad 3\cdot\lambda_{0}+\cdots+3\cdot\lambda_{\lfloor\frac{D+1}{3}\rfloor-1}+(D+2-3\cdot\left\lfloor\frac{D+1}{3}\right\rfloor)\cdot\lambda_{\lfloor\frac{D+1}{3}\rfloor}\geq 1,$ (7)
$\kappa=1:$ $\quad 3\cdot\lambda_{0}+\cdots+(D+5-3\cdot\lfloor\frac{D+1}{3}\rfloor)\cdot\lambda_{\lfloor\frac{D+1}{3}\rfloor-1}+\geq 1,$
$\vdots$
$\kappa=\lfloor\frac{D+1}{3}\rfloor:$ $\quad(D+2)\cdot\lambda_{0}\geq 1.$
Each level $\kappa$ in Eq. (7) is weaker than the previous $\kappa-1$, but also it does not require the knowledge of eigenvalue $\lambda_{\kappa}$. Notice that it also allows for some intermediate eigenvalues to be unknown. By considering the two smallest eigenvalues, we derive the weaker condition $(D-1)\lambda_{1}+3\lambda_{0}\geq 1$, which already includes the $2D$ vertices of both simplexes despite being suboptimal. The whole polytope of AS states is enclosed by all the possible permutations on the eigenvalues of inequality (6), but just as with the case of the simplexes in Figure 1, one linear inequality is enough for ordered eigenvalues (see Appendix A, B).
3.2 Improving Gurvits-Barnum criterion for arbitrary dimensions
Let us observe that the condition $\lambda_{\max}(\rho)\leq 1/(N\cdot M-1)$ is completely enclosed inside the Gurvits-Barnum ball. Thus, we focus on the set of states detected by the convex hull (CH) of the ball (Theorem 1) and the condition $\lambda_{\min}(\rho)\geq 1/(N\cdot M+2)$ (Theorem 3). This approach improves existing separability criteria for general dimension $D=N\cdot M$. Notably, the Hilbert-Schmidt distance between the vertices of the simplex and the maximally mixed state tends towards twice the radius of the ball as $N\cdot M\rightarrow\infty$. Figure 2 provides a 2D-sketch of the geometry of such a convex hull and how different regions are detected.
Figure 2: Detected absolutely separable (AS) states. Left panel: Schematic 2D representation of the convex hull (orange) of the Gurvits-Barnum ball (yellow) and our simplex (blue) in barycentric coordinates (c.f. Figure 1). Right panel: Purity $\mathrm{Tr}(\rho^{2})$ and minimal eigenvalue $\lambda_{\min}(\rho)$ of the detected states for a $2$-qutrit system. We distinguish the states detected as AS from Gurvits-Barnum ball (yellow region), the ones detected by the inverse of the reduction map (blue area), by both criteria (green region) and by its CH (orange area under the dotted curve). The points in the striped region do not correspond to physical states.
While an analytical expression for the CH (Figure 2) in higher dimensions is challenging to obtain, it is possible to state the problem as a disciplined convex program (DCP), as formalized in the next proposition.
Proposition 2.
Let $\rho$ be a bipartite state acting in a Hilbert space $\mathbb{C}^{N}\otimes\mathbb{C}^{M}$ of global dimension $D=N\cdot M$. If the problem
$\begin{array}[]{crl}\min_{\{\varrho_{1}\succeq 0,\varrho_{2}\succeq 0\}}&0&\\ \mbox{s.t.}&\rho&=\varrho_{1}+\varrho_{2}\\ &\lambda_{\min}(\varrho_{1})&\geq\frac{\mathrm{Tr}(\varrho_{1})}{D+2}\\ &\mathrm{Tr}(\varrho^{2}_{2})&\leq\frac{\mathrm{Tr}(\varrho_{2})^{2}}{D-1}\end{array}$ (8)
is feasible, then $\rho$ is absolutely separable (AS).
Proof.
If the problem is feasible, it implies that the algorithm indeed succeeded in finding $\{\varrho_{1},\varrho_{2}\}$ fulfilling the constraints. Hence, both states $\{\varrho_{1},\varrho_{2}\}$ are certified AS from Theorem 1 or Theorem 3, respectively. Finally, by convexity, their sum $\rho$ is also AS. ∎
3.3 Unifying separability criteria into a single convex program
Here, we outline how to generalize the disciplined convex program from Proposition 2 into potentially stronger sufficient conditions for separability. Although such criteria, as already discussed, arise from a plethora of different techniques, it is interesting to notice that all of them can be ultimately expressed as the collection of inequalities, $\{f_{i}(\rho)\leq 0\}$, where the functions $\{f_{i}\}$ are convex in the state $\rho$. Let us designate by $\mathcal{R}_{i}=\{\rho\succeq 0\mbox{ s.t. }f_{i}(\rho)\leq 0\}$ the set of (unnormalized) states detected by each criterion $i$. From the curvature of $\{f_{i}\}$, the sets $\{\mathcal{R}_{i}\}$ are convex as well. However, such sets, stemming from the difference conditions, may be incomparable. How can we build a new condition including them all? The following theorem offers a solution to this question and generalizes Proposition 2.
Theorem 4.
Let $\rho$ be a quantum state, and $\{f_{i}(\rho)\leq 0\}$, with $\{f_{i}\}$ a set of convex functions leading to incomparable sets of states, each of them representing a sufficient conditions for separability. If the convex program
$\begin{array}[]{crl}\min_{\{\varrho_{i}\succeq 0\}}&0&\\ \mbox{s.t.}&\rho&=\sum_{i}\varrho_{i}\\ &\{f_{i}(\varrho_{i})&\leq 0\}\\ \end{array}$ (9)
is feasible, then, $\rho$ is separable.
Proof.
The proof is analogous to the one of Proposition 2, where the conditions for AS are introduced as generic convex constraints. ∎
By construction, the previous theorem implicitly tests membership of the state $\rho$ in the convex hull [33] of the union of the detected sets $\bigcup_{i}\mathcal{R}_{i}:=\mathcal{R}$, in other words, it offers an answer to the question whether $\rho$ is contained in $\mathrm{CH}(\mathcal{R})$. Consequently, Theorem 4 potentially enables the certification of states $\rho$ that are not detected by any the criteria $\{f_{i}(\rho)\leq 0\}$ alone, but that nonetheless can be decomposed into states that are detected by such criteria. In turn, convex problems of the structure Eq. (9), are not only mathematically well-posed [34, 35], but for some instances they may admit polynomial-time solutions. In this regard, there is a number of available solvers tailored to these problems [36], as well as user-friendly parsers [37].
4 Multipartite absolute separability
In this section, we extend some of the previous results to the multipartite setting. A state $\rho\in\mathcal{B}(\mathbb{C}^{N}\otimes\mathbb{C}^{M}\otimes\dots\otimes\mathbb{C}^{T})$ is (fully) separable if it admits a convex decomposition onto product states
$\rho=\sum_{i}p_{i}\outerproduct{e_{i},f_{i},\dots,g_{i}}{e_{i},f_{i},\dots,g_{i}},$ (10)
[…]
$\sigma_{N-k}\hskip 4.33382pt\geq\hskip 4.33382pt\frac{d^{k}\cdot\mathds{1}}{d^{N}+2}\hskip 4.33382pt\geq\hskip 4.33382pt\frac{\mathds{1}}{d^{N-k}+2}.$ (11)
However, we recall that certifying separability with respect to all bipartitions in the state and its successive reduced matrices does not imply full separability [38, 39]. That is to say, certain multipartite states are entangled despite being separable with respect to its bipartitions. A paradigmatic example of such states is constructed from unextendible product bases (UPB) [40]. Biseparable entangled states also exist and have been characterized in the symmetric sector [41, 42].
The previous observation evidences the challenges to certify full separability as a much more stringent condition than biseparability. We begin by extending the concept of the Gurvits-Barnum ball [Eq. (1)] to the multipartite setting [43]. This norm-based condition represents one of the few known sufficient criteria for full separability, as we discuss below.
Lemma 5.
[43] Let $\rho$ be a normalized $N$-qudit state acting on a Hilbert space of total dimension $D=d^{N}$. If
[…]
with $A=2^{2-N}$, then $\rho$ is absolutely fully separable.
For the specific case of $N$ qubits (i.e., $d=2$), this bound can be improved to $A=\frac{\beta\cdot 2^{N}}{\beta+3^{N}}$, where $\beta=54/17$ [17].
In the same spirit as Theorem 3, Lemma 5 allows us to formalize the following sufficient condition for AS, stemming from the reduction-like map Eq. (2).
Theorem 6.
Let $\rho$ be an $N-$qudit state of total dimension $D=d^{N}$ and let $\Lambda_{\alpha}(\rho)=\mathrm{Tr}(\rho)\mathds{1}+\alpha\rho$ be a family of reduction-like maps. If
$\alpha_{-}^{*}\leq\alpha\leq\alpha_{+}^{*},$ (13)
then $\sigma=\Lambda_{\alpha}(\rho)$ is fully separable, where $\alpha_{\pm}^{*}=\frac{A\pm\sqrt{A\cdot(D-1)\cdot(D-A)}}{D-A-1}$ and $A=2^{2-N}$.
For the specific case $d=2$ we have $A=\frac{\beta\cdot 2^{N}}{\beta+3^{N}}$, being $\beta=54/17$.
Proof.
[…]
By construction, Theorem 6 does not detect separability beyond Lemma 5. However, as opposed to the condition given in Eq. (12), our criteria depend only on the minimal or maximal eigenvalue of the density matrix and not on its full spectrum. In the following, we specify the new condition.
Corollary 1.
Let $\sigma$ be an $N-$qudit state with minimal eigenvalue $\lambda_{\min}(\sigma)$ and maximal eigenvalue $\lambda_{\max}(\sigma)$. If
$\lambda_{\min}(\sigma)\geq\frac{1}{d^{N}+\alpha_{+}}\mbox{ or }\lambda_{\max}(\sigma)\leq\frac{1}{d^{N}-\alpha_{-}},$ (14)
where $\alpha_{\pm}$ are defined in Theorem 6, then $\sigma$ is fully separable.
Proof.
The proof in analogous to that of Theorem 3. ∎
Surprisingly, in contrast to its bipartite counterpart, the scalable condition Eq. (5) is not the tightest of this form. Already for $N=2$ it is weaker than Eq. (1). In fact, to date, the tight value of $A$ remains unknown. However, in [17], by exploring constructions of separable states at the border of entanglement, the authors report the upper bound $\beta=4$. Remarkably, such bound is also of order $\mathcal{O}(1)$ in the system’s size $N$, which implies that the gap between lower and upper bounds is closed up to a constant factor.
The existence of multipartite states fulfilling Theorem 3 but remaining unconsidered by Lemma 5 introduces the notion of a subtle form of entanglement: states that are AS and AP with respect to any bipartition and with respect to any bipartition of any of the reduced matrices, but yet not fully separable. In light of the intricacies of multipartite AS, in the next section we focus on the symmetric subspace, in which the characterization of these features is more manageable.
4.1 Symmetric absolute separability and absolute PPT
The notions of AS and AP introduced earlier may be immediately restricted to the symmetric subspace, which we shall denote here as $S(\mathcal{H})$. In a generic $N-$qudit system, $\mathcal{H}=(\mathbb{C}^{d})^{\otimes N}$, this subspace has dimension $\binom{N+d-1}{d-1}$. Specifically, we define
Definition 3.
A symmetric state $\rho$ acting on $S(\mathcal{H})$ is symmetric absolutely separable (or SAS) if and only if, given any unitary matrix acting on the symmetric subspace $U_{S}$, $\rho^{\prime}=U_{S}\rho_{S}U_{S}^{{\dagger}}$ is (fully) separable.
In addition, we introduce the weaker concept of symmetric absolutely PPT
Definition 4.
A symmetric state $\rho$ acting on $S(\mathcal{H})$ is symmetric absolutely PPT (or SAP) if and only if, given any unitary matrix acting on the symmetric subspace $U_{S}$, $\rho^{\prime}=U_{S}\rho_{S}U_{S}^{{\dagger}}$ is PPT with respect to any bipartition.
Let us start with the simplest case corresponding to the symmetric subspace of a $2-$qubit system. Following the approach from Theorem 2, that is, considering the inverse of the reduction map (3), the corresponding bounds on $\alpha\in[-3/4,1]$ that allow to certify SAS have been already derived in the general context of separability [11]. For SAS, such bounds can be reformulated as:
Theorem 7.
Let $\rho\in\mathcal{S}(\mathbb{C}^{2}\otimes\mathbb{C}^{2})$ be a symmetric state of a $2-$qubit system with minimal and maximal eigenvalues $\lambda_{\min}(\rho)$ and $\lambda_{\max}(\rho)$ respectively. If
$\lambda_{\min}(\rho)\geq\frac{1}{4}\quad\text{or}\quad\lambda_{\max}(\rho)\leq\frac{4}{9},$ (15)
then, $\rho$ is symmetric absolutely separable.
Proof.
The proof is analogous to that of Theorem 3, and uses the derivations on the inverse of the reduction map introduced by some of us in [11]. ∎
From the two simplexes described by the minimal and maximal eigenvalue conditions, we derive the following sufficient condition for SAS, as done similarly in Proposition 1.
Proposition 3.
Let $\rho\in\mathcal{S}(\mathbb{C}^{2}\otimes\mathbb{C}^{2})$ be a symmetric state of a $2-$qubit system with $\lambda_{0}\leq\lambda_{1}\leq\lambda_{2}$ its corresponding eigenvalues in non-decreasing order. If
$7\lambda_{0}+5\lambda_{1}\geq 3,$ (16)
then $\rho$ is symmetric absolutely separable.
Proof.
In Appendix A we show how to compute the linear inequality and verify how it is the most external one enclosing the ordered CH under ordered eigenvalues. ∎
However, as appreciated in Figure 3, even though the simplexes are tight, the CH of both of them does not attain the criterion $\sqrt{\lambda_{1}}+\sqrt{\lambda_{0}}\geq 1$ (see [19, 44]), which fully characterizes SAS in the simple $2-$qubit scenario.
Figure 3: Left panel: Comparison of the states detected for a symmetric $2-$qubit system with the convex hull (green line) of the sufficient conditions through the inverse of the linear map with $\alpha=1$ (dashed red line) and $\alpha=-3/4$ (dotted orange line) with the complete characterization $\sqrt{\lambda_{0}}+\sqrt{\lambda_{1}}\geq 0$ (black solid line). Right panel: Comparison of the states detected for a symmetric $3-$qubit system with the convex hull (green line) of the sufficient conditions through the inverse of the linear map with $\alpha=2/3$ (dashed red line) and $\alpha=-2/3$ (dotted orange line) with the best known criterion [45] (black sphere). The representation shows a $2D$ cut of the non-increasingly ordered region of eigenvalues $\lambda_{0}\leq\lambda_{1}\leq\lambda_{2}\leq\lambda_{3}$.
Next, for the second-simplest case of a symmetric $3-$qubit system, the bounds of $\alpha\in[-1/3,2/3]$ have also been derived analytically [11], allowing us to prove SAS from the inverse of the reduction map $\Lambda_{\alpha}^{-1}(\rho)$ (3). Nevertheless, numerical evidence [44] shows that the state with eigenvalues $(1,3,3,3)/10=(1-(N)/(N+1+\alpha),1/(N+1+\alpha),1/(N+1+\alpha),1/(N+1+\alpha))$ is still SAS. From this discussion, we infer that the tighter lower bound $\alpha=-2/3$ can be used to prove SAS.
Theorem 8.
Let $\rho\in\mathcal{S}(\mathbb{C}^{2}\otimes\mathbb{C}^{2}\otimes\mathbb{C}^{2})$ be a symmetric state of a $3-$qubit system with minimal and maximal eigenvalues $\lambda_{\min}(\rho)$ and $\lambda_{\max}(\rho)$ respectively. If
$\lambda_{\min}(\rho)\geq\frac{3}{14}\quad\text{or}\quad\lambda_{\max}(\rho)\leq\frac{3}{10},$ (17)
then, $\rho$ is symmetric absolutely separable.
Proof.
The proof is analogous to the one in Theorem 3. The positive bound of $\alpha=2/3$ has been derived in Ref [11], whereas the negative bound of $\alpha=-2/3$ has been obtained numerically. It is enough to prove it for pure states $\rho=\outerproduct{\Psi}{\Psi}$ |. For $3-$qubits, |$\ket{\Psi}$ has $4$ complex coefficients in the symmetric (Dicke) basis |$\ket{\Psi}=\sum_{i}x_{i}\ket{D_{i}^{(3)}}$. It is possible to minimize for all $\alpha\in[-1,0]$ over the coefficients $\{x_{i}\}$ the smallest eigenvalue of $\Lambda_{\alpha}^{T_{2:4}}(\rho)$ and check that $\alpha=-2/3$ is the smallest value of $\alpha$ that cannot lead to a negative partial transpose. The value is also considered extreme in [44], and we have extended their numerical calculations of the maximal possible negativity for all the states with eigenvalues $(1,3,3,3)/10$ with over 2 million random unitary matrices. ∎
Remarkably, by characterizing the CH of both simplexes, we provide a stronger sufficient condition for SAS from few eigenvalues.
Proposition 4.
Let $\rho\in\mathcal{S}(\mathbb{C}^{2}\otimes\mathbb{C}^{2}\otimes\mathbb{C}^{2})$ be a symmetric state of a $3-$qubit system and $\lambda_{0}\leq\lambda_{1}\leq\lambda_{2}\leq\lambda_{3}$ its corresponding eigenvalues in non-decreasing order. If
$6\lambda_{0}+6\lambda_{1}+2\lambda_{2}\geq 3,$ (18)
then $\rho$ is symmetric absolutely separable.
Proof.
The proof is analogous to the one of Proposition 3. For the analytical lower bound of $\alpha\in[-1/3,2/3]$ derived in [11], the linear CH obtained is given by the expression $5\cdot\lambda_{1}+9\cdot\lambda_{0}\geq 3$. ∎
Interestingly, for $N=3$ symmetric qubits, some AS states lie outside the region given by Eq. (18), as recently reported [44]. Nevertheless, our analytical linear expression includes all known, to date, analytical criteria [45], and provides the best existing bound for SAS for this specific set of states. In Figure 3, we display the known conditions that detect SAS states, together with those derived in this work. As it can be seen, the volume of SAS-states detected with our conditions represents a notable improvement.
Regarding SAP (symmetric absolute PPT), let us recall that in the symmetric subspaces $S(\mathbb{C}^{d}\otimes\mathbb{C}^{d})$ for $d=2,3$, PPT is a sufficient condition for separability, and thus SAS is equivalent to SAP in such cases. In what follows, we derive new bounds to certify SAP for the general $N-$qudit symmetric spaces with respect to any bipartitation. The specific case of SAP for bipartite qudit systems, i.e. $N=2$, has been completely characterized by a series of matrix inequalities on the eigenvalues of such systems [19]. However, the number of such matrices grows fast with the local dimension $d$. Thus, the conditions we present here are much more practical.
Theorem 9.
[…]
$-\binom{N}{\lfloor N/2\rfloor}^{-1}=-\lambda_{min}\leq\alpha\leq 2\lambda_{min}=2\cdot\binom{N}{\lfloor N/2\rfloor}^{-1},$
where $\lambda_{min}>0$ is the minimal eigenvalue of $\mathds{1}_{S}^{T_{A}}$ and the partition $A$ includes $\lfloor N/2\rfloor$ qudits. Then, $\rho_{S}\geq 0\rightarrow\Lambda_{\alpha}(\rho_{S})=\sigma_{S}$ is SAP.
Similarly, if $\sigma_{S}\geq 0$ and $\Lambda_{\alpha}^{-1}(\sigma_{S})=\rho_{S}\geq 0$, then $\sigma_{S}$ is SAP.
Proof.
[…]
Some interesting remarks stem from Theorem 9. First, symmetric states of $N-$qudits that are absolutely PPT in the partition $11:N-$ 1 can still be NPT with respect to some larger partition. Since the set of SAP states with respect to the largest bipartition is the smallest one, and it is contained in the sets of SAP of smaller partitions, we will consider it as the one defining SAP 4. It is also necessary to mention that we have numerical evidence that the upper bound of $\alpha_{+}=2\cdot\binom{N}{\lfloor N/2\rfloor}^{-1}$ is tight for the SAP set, since increasing such value by even a $0.1\%$ leads to NPT states under certain unitary transformations. Nevertheless, the same behavior has not been found for the negative lower bound on $\alpha_{-}$. As an example, one can check that for $N=2,3$ qubits, the lowest bounds are given by $\alpha_{-}=-3/4,-2/3$ respectively, instead of the values $\alpha_{-}=-1/2,-1/3$ given by Theorem 9.
The previous theorem leads to the following corollary as a sufficient condition for SAP in $N-$qudit systems.
Corollary 2.
Let $\rho\in\mathcal{S}((\mathbb{C}^{d})^{\otimes N})$ be a symmetric state of an $N-$qudit system of dimension $D=d^{N}$ with minimal and maximal eigenvalue $\lambda_{\min}(\rho)$ and $\lambda_{\max}(\rho)$ respectively. If
$\lambda_{\min}(\rho)\geq\frac{1}{d^{N}+2\cdot\binom{N}{\lfloor N/2\rfloor}^{-1}}\text{ or }\lambda_{\max}(\rho)\leq\frac{1}{d^{N}-\binom{N}{\lfloor N/2\rfloor}^{-1}},$ (19)
then $\rho$ is symmetric absolutely PPT.
Proof.
The proof is analogous to that of Theorem 3. ∎
For the specific case of N-qubit systems, a general linear condition on the eigenvalues of a given state $\rho$ can be provided to ensure it is SAP, in a complete analogy as it was demonstrated in Proposition 1. Moreover, the same procedure can be generalized for N- qudit systems, $S((\mathbb{C}^{d}))^{\otimes N})$, even though we do not explicitly describe it in this work.
Proposition 5.
Let $\rho\in\mathcal{S}((\mathbb{C}^{2})^{\otimes N})$ be a symmetric state of $N$ qubits and $\{\lambda_{i}\}_{i=0}^{N}$ its eigenvalues in non-decreasing order $\{\lambda_{i}\leq\lambda_{i+1}\}_{i=0,1,..,N}$. If $\rho$ is contained in the CH of the conditions presented in Corollary 2, then
$3\cdot\binom{N}{\lfloor N/2\rfloor}\sum_{i=0}^{\lfloor\frac{N+1}{3}\rfloor-1}\lambda_{i}+\left[\binom{N}{\lfloor N/2\rfloor}\cdot(N+1-3\cdot\lfloor\frac{N+1}{3}\rfloor)+2\right]\cdot\lambda_{\lfloor\frac{N+1}{3}\rfloor}\geq\binom{N}{\lfloor N/2\rfloor}.$ (20)
Proof.
The proof is analogous to that of Proposition 1. ∎
When it comes to comparing our criterion in Eq. (20) with existing SAS criteria [22, 44, 45], we find the extremal states detected by our conditions not captured by any of these. Thus, our condition provides a broader set of states detected. Nevertheless, our proof considers only SAP, for which we have not found analytical expressions in the literature. The sets of (S)AS and (S)AP being equivalent remains an open problem to possibly consider our criteria for separability or to disprove the conjectured equality of both sets [24].
The negative bound on $\alpha_{-}$ in Theorem 9, as already discussed, is not tight in general, and we foresee a dependence on $d$ for it. From [19], it is possible to compute the exact bound for the symmetric bipartite case, when $N=2$ and the local dimension $d$ is arbitrary. For the specific case of a symmetric $2-$qutrit system, we find the following results, which again only depend on a single eigenvalue of the density matrix.
Theorem 10.
Let $\rho\in\mathcal{S}(\mathbb{C}^{3}\otimes\mathbb{C}^{3})$ be a symmetric state of a $2-$qutrit system with minimal and maximal eigenvalues $\lambda_{\min}(\rho)$ and $\lambda_{\max}(\rho)$ respectively. If
$\lambda_{\min}(\rho)\geq\frac{1}{7}\quad\text{or}\quad\lambda_{\max}(\rho)\leq\frac{3}{16},$ (21)
then, $\rho$ is symmetric absolutely PPT.
Proof.
The positive and negative bounds on $\alpha\in[-2/3,1]$ can be computed by substituting the vector of eigenvalues $\left(1-(6-1)/(6+\alpha_{\pm},1/(6+\alpha_{\pm}),\cdots,1/(6+\alpha_{\pm}))\right)$ on Corollary 3 of [19]. The remainder of the proof is analogous to the one of Theorem 3. ∎
From the two simplexes given by the minimal and maximal eigenvalue criteria, we derive the following sufficient condition, as done previously for similar cases.
Proposition 6.
Let $\rho\in\mathcal{S}(\mathbb{C}^{3}\otimes\mathbb{C}^{3})$ be a symmetric state of a $2-$qutrit system with $\{\lambda_{i}\leq\lambda_{i+1}\}_{i=0,\cdots,5}$ its corresponding eigenvalues in non-decreasing order. If
$5\lambda_{0}+5\lambda_{1}+4\lambda_{2}\geq 2,$ (22)
then $\rho$ is symmetric absolutely PPT.
Proof.
The proof is analogous to that of Proposition 3. ∎
For higher dimensional bipartite systems $d>3$, one may recover a tight value of negative $\alpha$ by relaxing the LMIs presented in [19]. However, the number of such LMIs to consider prohibits a scalable characterization beyond $d=4$. For this last case, we derive the bound $\alpha_{-}=-0.625$ from their Conjecture 1. Strong numerical evidence suggests that such bound is indeed tight. Thus providing the following criteria for two ququart symmetric states.
Theorem 11.
Let $\rho\in\mathcal{S}(\mathbb{C}^{4}\otimes\mathbb{C}^{4})$ be a symmetric state of a $2-$ququart system with minimal and maximal eigenvalues $\lambda_{\min}(\rho)$ and $\lambda_{\max}(\rho)$ respectively. If
$\lambda_{\min}(\rho)\geq\frac{1}{11}\quad\text{or}\quad\lambda_{\max}(\rho)\leq\frac{8}{75},$ (23)
then, $\rho$ is symmetric absolutely PPT.
Proof.
The positive and negative bounds on $\alpha\in[-5/8,1]$ can be computed by substituting the vector of eigenvalues $\left(1-(10-1)/(10+\alpha_{\pm},1/(10+\alpha_{\pm}),\cdots,1/(10+\alpha_{\pm}))\right)$ on Conjecture 1 of [19]. We have gathered numerical evidence that the value of $\alpha_{-}$ is, in fact, tight, supporting the conjecture. The rest of the proof is analogous to the one of Theorem 3. ∎
From the two simplexes given by the minimal and maximal eigenvalue conditions, we derive the following sufficient condition, as done previously for similar cases.
Proposition 7.
Let $\rho\in\mathcal{S}(\mathbb{C}^{4}\otimes\mathbb{C}^{4})$ be a symmetric state of a $2-$ququart system with $\{\lambda_{i}\leq\lambda_{i+1}\}_{i=0,\cdots,9}$ its corresponding eigenvalues in non-decreasing order. If
$3\lambda_{4}+13\cdot\sum_{i=0}^{3}\lambda_{i}\geq 5,$ (24)
then $\rho$ is symmetric absolutely PPT.
Proof.
The proof is analogous to that of Proposition 3. ∎
To close the section, we summarize the main criteria established here for SAS. We first provide tight SAS criteria for two and three qubits. Then, we examine the twin problem of SAP and present new non-tight criteria for arbitrary dimensions and tight ones for two qutrits and ququarts systems. For these, we have strong numerical evidence for the tightness of our bounds. Finally, it would be beneficial to develop scalable criteria for SAP and study its interplay with SAS, which we leave for subsequent investigations.
5 Conclusions
In this work we have presented a new method to tackle the problem of absolute separability (AS) and absolute PPT (AP) based on inverting non completely positive maps. The AS (AP) sets are defined as those formed by the states whose separability (positivity under partial transposition) is preserved under any global unitary transformation. These sets are convex and contain the maximally mixed state as well as states sufficiently close to it. Due to this invariant condition, AS states are completely determined from their spectrum.
Our method provides novel simple analytical sufficient conditions for AS in arbitrary dimensions based on the knowledge of very few or even one eigenvalue of the spectrum. These analytical conditions correspond to convex simplexes that contain extremal points of the AS set, improving presently known conditions for AS. Moreover, we have demonstrated that AS conditions can be cast as convex functions of the form $f(\rho)\leq 0$. This fact allows us to unify all known AS conditions into a single and stronger one using convex programming. Such a technique can be generalized to any convex set to improve its characterization, which is normally a very hard task. Our method also yields new conditions on the problem of absolute separability in the multipartite scenario, as well as in the case of the symmetric subspace. Finally, we provide some new criteria on the twin problem of AP.
Let us conclude that the criteria presented in this work rely on the knowledge of the lowest eigenvalues of the state. In contrast, the majority of setups aimed at the evaluation of eigenvalues focus on the maximal one [48, 49]. However, our results motivate the study of the experimental setups that lower-bound the spectrum of a given state $\rho$, prompting new theoretical and experimental avenues on the problem of separability.
Acknowledgements.– We thank A.Diebra and J.Romero-Pallejà for insightful comments and discussions with J. Ahiable, K.N.B. Teja, J.Tura and A. Winter.
We acknowledge financial support from Spanish MICIN (projects: PID2022:141283NB-I00;139099NBI00) with the support of FEDER funds, the Spanish Goverment with funding from European Union NextGenerationEU (PRTR-C17.I1), the Generalitat de Catalunya, the Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project -Quantum Spain Project- through the Recovery, Transformation and Resilience Plan NextGeneration EU within the framework of the Digital Spain 2026 Agenda. We also acknowledge support from: Europea Research Council AdG NOQIA; MCIN/AEI (PGC2018-0910.13039/501100011033, CEX2019-000910-S/10.13039/501100011033), EU-QUANTera Projects: MAQS PCI2019-111828-2, DYNAMITE PCI2022-132919, ExTRaQT-PCI2022-132965. Support from Fundació Cellex; Fundació Mir-Puig; Generalitat de Catalunya (European Social Fund FEDER and CERCA program, AGAUR Grant No. 2021 SGR 01452, QuantumCAT  U16-011424, co-funded by ERDF Operational Program of Catalonia 2014-2020); Barcelona Supercomputing Center MareNostrum (FI-2023-3-0024) funded by the European Union are acknowledged. This project has received funding from EU Horizon Europe Program ORIZON-CL4-2022-QUANTUM-02-SGA PASQuanS2.1, 101113690; FET-OPEN OPTOlogic, Grant No 899794, NeQSTGrant-01080086. ICFO Internal “QuantumGaudi" project. Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Derivation of a single linear inequality from the convex hull of two simplexes
In this appendix, we present the detailed procedure for deriving the inequalities described in Eqs. (6), (16), (18), (20), (22) and (24) which arise from the geometric structure of the two simplexes in the eigenvalue space, given by positivity of the inverse of the reduction map $\Lambda_{\alpha}^{-1}(\rho)\geq 0$ (3). For the low-dimensional specific cases, we also describe the procedure to check they become an AS condition, while giving numerical evidence for arbitrary dimensions that Eqs. 6 and 20 to also hold as sufficient conditions for AS, besides containing the whole CH of the two linearly independent simplexes (proven in Appendix B).
In the context of absolute separability (AS) of systems of total dimension $D$, sufficient conditions are inherently dependent on the spectrum $\bm{\lambda}$ of the quantum state $\rho$. Consequently, it is often more practical to work within the space of eigenvalues $\{\lambda_{i}\}$ rather than over the entire space of physical states.
From the positivity of the inverse of the reduction map (3), we derive $D$ inequalities for the $D$ eigenvalues of the state $\rho$ for each value of $\alpha$. Theorem 3 provides a simplified form of these inequalities, given a specific ordering of the eigenvalues. The $D$ conditions enclose a regular $(D-1)$-dimensional simplex, which center is in the maximally mixed state (MMS). The vertices of the simplex are given by all the different permutations of the vector $(1-(D-1)/(D+\alpha),1/(D+\alpha),\cdots,1/(D+\alpha))$.
The unit trace constraint allows us to eliminate a single variable, $\sum_{i}\lambda_{i}:=\tilde{\lambda}_{0}=1$. Next, the orthogonal complement of the vector $(1,\cdots,1)$ is used to complete a basis of the $D-1$ subspace describing the barycentric coordinates (see Figure 1 of the main text).
The convex hull (CH) formed by both simplexes represents a polytope whose vertices are the union of the vertex sets from each simplex. While the resulting polytope is not regular, it exhibits symmetry under eigenvalue permutations, reflecting the identical ordering possibilities of the eigenvalues. In computational geometry, the $H-$representation (half-space representation) of a polytope involves describing it as an intersection of half-spaces. These half-spaces can be represented by linear inequalities of the form $A\bm{\lambda}\leq\mathbf{b}$, being $A$ a matrix, but since the desired object is invariant under permutations, the matrix $A$ will be in this case a matrix of $D!$ identical rows (the possible $D!$ orderings of the eigenvalues [50]).
For low dimensional systems, it is possible to compute the set of linear inequalities analytically, by computing the hyperplanes given by each possible set of $D-1$ points, and considering the most external ones as the ones containing the vertices of the biggest dual polytope. Given the difficulty of extending the calculations for all $D$, a numerical approach has been used to compute the equations of the facets of the polyhedron. In specific, making use of the python library pypoman - Polyhedron manipulation in Python [51]. The direct output of the package is the matrix $A$ and the vector $\mathbf{b}$ for a given set of points (the vertex of both simplexes).
In $D-1$-dimensional space, these inequalities depend on the orientation of the basis orthogonal to the vector $(1,\cdots,1)$, not only on the dimension of the Hilbert space. To generalize these inequalities across different dimensional systems, we reexpress them in the full $D$-dimensional eigenvalue space. This allows for a more straightforward comparison between different quantum systems, and the resulting linear matrix inequalities take the form $A\bm{\lambda}\leq\mathbf{b}$, where $\mathbf{b}$ is a constant vector, and $A$ is a matrix of permuted vectors $\mathbf{c}=(p,\cdots,p,r,-q,\cdots,-q)$. Here, the multiplicities of $p$, $q$, and $r$ are functions of the system’s dimension, ensuring that the inequalities remain invariant under permutations. Being $m(p)$ the multiplicity of $p$ in the given vector, the size of the matrix is given by $D\times\binom{D}{\max{m(p),m(q)}}\leq D!$. Where $p,r,q,s$ can be expressed as a function of the total dimension of the system.
It is necessary to find the expression that corresponds to the ordered eigenvalues, i.e., the one that encloses the specific region in the polytope in non-decreasing order ($\mathbf{c}$). Reapplying the normalization condition $\sum_{i}\lambda_{i}=1$, one can eliminate the dependence on all the eigenvalues whose coefficient would be $p$ in $\mathbf{c}\cdot\bm{\lambda}^{T}$.
It is important to note that these expressions have not been derived in a general manner, but instead from a bottom-up approach, where we have computed the first low-dimensional cases and generalized the dependence of $p,q,r,s$ on the total dimension of the system, checking the validity for some high dimensional cases.
In general, it is possible to find in a case by case basis the $D-1$ points that generate the hyperplane, as well as to check that it encloses the bigger dual polytope. By considering the intersection of the external hyperplanes with the ordering, it is also possible to see that one single inequality is enough to establish the criteria. Nevertheless, for high dimensions, this procedure becomes intractable even though the python library computes the most external ones by construction.
An example of some gathered numerical evidence can be done by uniformly sampling random states. Then, the disciplined convex program based on Theorem 4:
$\begin{array}[]{crl}\min_{\{\varrho_{1},\varrho_{2}\succeq 0\}}&0&\\ \mbox{s.t.}&\rho&=\varrho_{1}+\varrho_{2}\\ &\lambda_{\min}(\varrho_{1})&\geq\frac{\mathrm{Tr}(\varrho_{1})}{D+\alpha_{+}}\\ &\lambda_{\max}(\varrho_{2})&\leq\frac{\mathrm{Tr}(\varrho_{2})}{D-|\alpha_{-}|}\\ \end{array},$ (25)
being $\alpha_{\pm}$ the extreme values of $\alpha$, identifies the same sets of states as those found through the linear inequalities derived from the eigenvalues.
Appendix B Proof that the convex hull satisfies Proposition 1
After obtaining numerically the convex hull inequality, as described in Appendix A, we proceed to verify that the convex hull of two simplexes satisfies such inequality. To this end, we revisit key concepts of majorization, as outlined in [52], along with the references within the text.
To begin with, if a vector $\mathbf{x}$ is majorized by $\mathbf{y}$, that is, $\mathbf{x}\prec\mathbf{y}$ then $\mathbf{x}$ is less spread than $\mathbf{y}$. In the context of the space of eigenvalues of quantum density matrices, we refer as $\mathbf{x}$ being more mixed than $\mathbf{y}$. Formally,
Definition 5.
Consider two $D$-dimensional vectors $\mathbf{x}$ and $\mathbf{y}$ ordered in a non-decreasing way ($x_{0}\leq\cdots\leq x_{D-1}$), we say that $\mathbf{x}\prec\mathbf{y}$ if
$\sum_{i=0}^{D-1}x_{i}=\sum_{i=0}^{D-1}y_{i}$ (26)
$\sum_{i=0}^{k}x_{i}\geq\sum_{i=0}^{k}y_{i},\quad k=0,\cdots,D-2$ (27)
We also present two lemmas, that will be of great utility.
Lemma 12.
A given vector $\mathbf{x}$ is majorized by another vector $\mathbf{y}$ if and only if $\mathbf{x}$ is in the convex hull of the $D!$ permutations of $\mathbf{y}$.
From the previous lemma, one can immediately infer that any point in the space of eigenvalues of a density matrix inside the simplexes given by Eq. (3) is majorized by any of the extreme points of the simplex, which are all different permutations of a single vertex.
Lemma 13.
[52, 53] If $\mathbf{a}\prec\mathbf{b}$, $\mathbf{u}\prec\mathbf{v}$ and if $\mathbf{b}$ and $\mathbf{v}$ are similarly ordered, then $\mathbf{a}+\mathbf{u}\prec\mathbf{b}+\mathbf{v}$.
Since tha majorization condition requiers the fulfillement of the normalization Eq. (26), the previous lemma implies that any point inside of the convex hull of the two simplexes is majorized by some points on the border of the convex hull (obtained as a linear combination of the vertex of the simplexes).
Next, we introduce the concept of Schur-convex functions as those non-increasing under majorization:
Definition 6.
A function $f:\mathbb{R}^{D}\rightarrow\mathbb{R}$ is defined Schur-convex if and only if for all $\mathbf{x},\mathbf{y}\in\mathbb{R}^{D}$ such that $\mathbf{x}$ is majorized by $\mathbf{y}$ one has that $f(\mathbf{x})\leq f(\mathbf{y})$.
The next lemma provides a necessary and sufficient conditions for a function to be Schur-convex.
Lemma 14.
A function $f(\mathbf{x})$ is Schur-convex if it fulfills the following conditions:
  • $f(\mathbf{x})$ must be symmetric under permutations of the arguments of $\mathbf{x}$
  • $(x_{i}-x_{j})\cdot\left(\frac{\partial f}{\partial x_{i}}-\frac{\partial f}{\partial x_{j}}\right)\geq 0$ for all $\mathbf{x}\in\mathbb{R}^{D}$.
We now verify that any vector contained in the convex hull of the two simplexes fulfills the proposed condition. First, given the eigenvalues of a density matrix in non-decreasing order $\{\lambda_{i}\leq\lambda_{i+1}\}_{i=0,\cdots,D-1}$ we define
$f(\bm{\lambda})=1-3\cdot\sum_{i=0}^{\lfloor\frac{D+1}{3}\rfloor-1}\lambda_{i}+\left(D+2-3\cdot\left\lfloor\frac{D+1}{3}\right\rfloor\right)\cdot\lambda_{\lfloor\frac{D+1}{3}\rfloor},$ (28)
from Eq. (6) such that our condition is expressed as $f(\bm{\lambda})\leq 0$ (c.f. Section 3.3). Here, the argument of the function is any vector of eigenvalues, which is not necessarily sorted. This function is clearly symmetric under permutations, as it accounts for the ordering of each component of the vector independently of its position. Furthermore, it can be verified from Lemma 14 that for natural values of $D,N\in\mathbb{N}$, the function $f(\bm{\lambda})$ is Schur-convex. On the other hand, we recall that the vertices of the AS simplexes are given by all the permutations of the vectors $\bm{\lambda}_{\pm}=\left(\frac{1}{D+\alpha_{\pm}},\cdots,\frac{1}{D+\alpha_{\pm}},1-\frac{D-1}{D+\alpha_{\pm}}\right)$, with $\alpha_{+}=2,\alpha_{-}=-1$ in the space of eigenvalues. Due to the convexity of the AS set and Schur-convexity of the function, we check the inequality $f(\bm{\lambda})\leq 0$ on the extremals $\bm{\lambda}_{\pm}$. In this regard, we confirm the tight value $f(\bm{\lambda}_{\pm})=0$. Thus, it follows from Schur-convexity that all the vectors contained in the convex hull of the points generated by the permutations of $\bm{\lambda}_{\pm}$ fulfill the inequality. Finally, the procedure outlined here can be extended to Eqs. (16), (18), (20), (22) and (24) of the main text.